What is the cost of debt capital given the following information?

Number of bonds = 1,000,000

Face value = 1,000

Current price = 1,100

Coupon rate = 9% semiannually

Years of maturity = 15 years




1
Expert's answer
2021-03-25T19:27:47-0400

The cost of debt of bonds is equal to the Yield to maturity of bonds.

"YTM=\frac{Coupon+(FV-PV)/n}{(FV+PV)/2}"

"Coupon = 0.09*2*1000=180"

"YTM=\frac{180+(1,000-1,100)/15}{(1,000+1,100)/2}=16,51" %